OpenStudy (anonymous):

find the value of k if 2x^2-kx+8=0

OpenStudy (anonymous):

a=2, b=-k, c=8

OpenStudy (anonymous):

discriminant b^2-4ac

OpenStudy (anonymous):

how do you get k?

OpenStudy (campbell_st):

well you would need some condition if you need to use the discriminant...like the roots are equal.... do discriminant = 0 roots are real, unequal and rational... etc... but looking at the problem k could be -8, -10, -17,

OpenStudy (anonymous):

please how did you arrive at -8, -10, -17

OpenStudy (campbell_st):

well look at the factors of 8.... they are -1 and -8 then -2 and -4 this is because the middle term is -k... so factoring (2x -1)(x - 8) k = -17 or (2x - 8)(x -1) k = -10 (2x -2)(x - 4) k = -10 or (2x - 4)(x -2) k = -8 thats providing that the quadratic can be factored... if not... then there are an infinite number of solutions... thats why you need some extra information

OpenStudy (anonymous):

really

OpenStudy (anonymous):

can you please try using the discriminant formula

OpenStudy (anonymous):

(-k)^2-4(2)(8)=0

OpenStudy (campbell_st):

well if you use the discriminant all you get is an expression \[k^2 - 4 \times 2 \times 8\] so \[k^2 - 64\] it can be factored... but that all...as its just an expression

OpenStudy (campbell_st):

so are you given the fact that the discriminant is equal to zero...?

OpenStudy (anonymous):

(-k)^2 shoud equal k

OpenStudy (anonymous):

why k^2

OpenStudy (campbell_st):

well (-k)^2 = k^2

OpenStudy (campbell_st):

its -k x -k = k^2 the negative times a negative.... thing

OpenStudy (anonymous):

oh ok